OpenStudy (anonymous):

2x – 3y = 12 -x – 3y = -6 find the points.

hero (hero):

Divide the second equation by -1, then add both equations together.

OpenStudy (anonymous):

If you subtract the two equations (this is the same as multiplying one of the equations by -1 and then adding them), the terms with y will cancel out 2x – 3y = 12 - ( -x – 3y = -6) ------------------ 3x = 18 x = 6 You can plug that back into on of the original equations to solve for y 2(6) - 3y = 12 -3y = 0 y = 0